OpenStudy (anonymous):

can someone show the steps in solving lne^4

OpenStudy (goformit100):

do you want the expansion of it ?

OpenStudy (anonymous):

yes please!

OpenStudy (tkhunny):

4 -- done. What really is wanted?

OpenStudy (anonymous):

@tkhunny, I think they were looking more for an explanation with principles instead of the answer; though your answer is completely correct! Well the principle of this is that: \[\log_a{(a^x)}=x\] Reason being is we know that \(\log_a{b^c}=c\times\log_a{b}\) So therefore, we simplify: \[\log_a{a^x}=x\log_a{a}=x(1)=x\] So therefore: \[\ln(e^4)=\log_e{e^4}=4\log_e{e}=4(1)=4\]

OpenStudy (anonymous):

thanks a ton, thats what I was looking for!